How do I work and solve the question?
The smallest angle in a triangle is 1/3 as large as the largest angle. The third angle is twice the smallest and. Find the three angles.

Answer by rfer(16322)

x+1/3x+2/3ex=180
2x=180
x=180/2
x=90
1/3x=30
2/3x=60